OpenStudy (noahbred):

Any help i will medal? solve -5/8c=20

OpenStudy (noahbred):

Soryy its not grad question i chose wrong one

OpenStudy (anonymous):

OK, take 8c to the other side which will become -5=160c. Know take the 160 to the other side and you get -5/160=c. Simplify it and you'll get -1/32=c

OpenStudy (noahbred):

Oh so pretty much the answer could just be -32 yah

OpenStudy (anonymous):

Actually \[-\frac{ 1 }{ 32 }\]

OpenStudy (anonymous):

Here is a website to help with such questions: https://mathway.com/

OpenStudy (noahbred):

but the ansers dont have that it only has this A. –32 B. –34 C. 32 D. 34

OpenStudy (anonymous):

Are you sure it's \[-\frac{ 5 }{ 8c }\] Don't you mean \[-\frac{ 8c }{ 5 }\]

OpenStudy (anonymous):

If it's the second one then it's -32

OpenStudy (noahbred):

 its this